what is the fraction notation for the ratio of 8640 to 14400
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)
Answer by stanbon(75887)

what is the fraction notation for the ratio of 8640 to 14400
----
Ans: 8640/14400 = 4320/7200 = 2160/3600 = 1080/1800
= 540/900 = 270/450 = 135/225 = 27/45 = 3/5
====================
Cheers,
Stan H.